OpenStudy (anonymous):

Please help me! I really want to understand this. How do you know if certain points are removable or non removable discontinuity without drawing a graph?

OpenStudy (anonymous):

well from limit

OpenStudy (anonymous):

A discontinuous function may be of three types

OpenStudy (anonymous):

I know but how can you tell from a given function. D: sometimes the function cant be graphed quickly ;_; like more complicated rationals..

OpenStudy (anonymous):

or maybe i cant do that? ;_;

OpenStudy (anonymous):

If the limit exists at a point, the discontinuity is removable.

OpenStudy (anonymous):

To check whether a limit exists, plug in points that are close to the point of discontinuity and check if it tends toward some value, or use limit properties to evaluate it.

OpenStudy (anonymous):

i still dont really understand because in one of my math problems, one point in it was removable and the other was non removable...

OpenStudy (anonymous):

when i tried making the equation work for both of them, the denominator would still be 0...

OpenStudy (anonymous):

http://i.imgur.com/Su0Jtqp.jpg

OpenStudy (anonymous):

Plug in 2.

OpenStudy (anonymous):

(1/2)(2)+1 = 1+1=2 3-(2) = 1

OpenStudy (anonymous):

They don't meet, so it's not continuous.